OpenStudy (anonymous):

It is given that AB=21, BC=14, and AD=54. Find the length of CD. ___._____.___.______.____ A B C D

OpenStudy (anonymous):

AD = AB+BC+CD 54 = 21+14+CD CD = ?

OpenStudy (anonymous):

19

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

find the supplement of a 7 degree angle